# Language Model Distillation: A Temporal Difference Imitation Learning Perspective

### v1

@@ Abstract
1 Introduction
The successes of large language models (LLMs) have brought transformative advancements across many NLP tasks, albeit often at the cost of substantial computational resources due to their massive size. Model distillation is a widely adopted approach to compressing these large, high-capacity models (teachers) into smaller, efficient models (students). This practice has become standard, ranging from proprietary ones [1] to open-source model families [2, 3, 4, 5, 6].
The seminal work of knowledge distillation (KD) proposed by Hinton et al. [7] established the foundation for distribution-matching distillation methods, which were later extended to sequence/auto-regressive models [8]. In distribution-matching methods, at each time step during text generation, a student model $\pi:\mathcal{X}\to\Delta(\mathcal{V})$ seeks to approximate the probability distribution, over an vocabulary $\mathcal{V}$, produced by the teacher $\pi^{\star}$, for example by minimizing a (forward) KL-divergence [9, 10, 11] $\min_{\pi}\mathbb{KL}[\pi^{\star}(x_{\leq t})\|\pi(x_{\leq t})]$, where $x_{\leq t}\in\mathcal{X}$ is a sequence of length $t$, and $\Delta(\Omega)$ denotes a probability simplex over space $\Omega$.
Notably, this sequential distribution matching can be interpreted as behavior cloning (BC) [12], within the imitation learning (IL) and inverse reinforcement learning (IRL) literature [13, 14, 15, 16, 17]. While there are nuanced differences between IL and IRL, we use them interchangeably.
@@ 1 Introduction
Casting distillation as an IL problem is not new per se; indeed, many existing distillation works [8, 18, 19, 20] perform BC. Before reviewing these works through the lens of IL, we highlight two key distinctions that help characterize the distillation setup: (i) offline vs. online: This refers to whether distillation is performed on a fixed, pre-collected dataset (offline) or involves continuously generating new data during training (online). (ii) off-policy vs. mixed-policy: A method is considered purely on-policy if it uses only student-generated data for training; otherwise, it is categorized as off-policy. However, purely on-policy distillation is rare in practice—most methods use a mixture of teacher and student data. We refer to this setup as mixed-policy, though it is technically off-policy in the RL literature.
The pioneer work of sequence-level KD (SeqKD) [8] is an offline off-policy BC. Follow up BC style works fall into two main categories: (i) online mixed-policy [18, 19, 21]; and (ii) offline mixed-policy [20]; potentially with different choice of divergence measure - forward/reverse KL [8, 19], Jensen-Shannon divergence [18], and general $f$-divergence [20].
Figure 1: We average the sorted token probabilities across $20$ sequences generated by Qwen-2.5 3B. Top $50$ tokens account for $96\%$ of the total mass, and the top $7$ tokens contribute $\geq 90\%$.
However, BC is known to suffer from compounding errors [22, 23], also referred as exposure bias [24] in the context of auto-regressive models. To address these issues, the IL literature offers a rich toolbox beyond BC, for example generative adversarial training [17], occupancy matching [25], and more. We refer readers to Osa et al. [26] for a comprehensive review. While not exhaustive, many modern IL methods [25, 17, 16] involve modeling an inverse reward function and then optimizing a policy with respect to it using RL. These RL algorithms are often based on temporal-difference (TD) learning [27], which estimates the long-term impact of an action (or token) given a state (or sub-sequence) to mitigate the compounding errors.
Distillation may potentially benefit from the broader IL toolbox—particularly from TD learning methods. Earlier work has already demonstrated the potential of such approaches in smaller auto-regressive models. For instance, Yu et al. [28] and Wu et al. [29] can be viewed as early applications of Ho and Ermon [17]. For modern LLMs, Jia [30] explored TD-style IL methods for distillation, through value moment matching [13, 16, 31] in particular. However, applying TD learning to LLMs is generally challenging due to their vast state-action space $\mathcal{V}^{T}$ [32, 33, 34].
@@
2 Preliminaries
MDP notations. We define the conventional MDP tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma)$ as follows. Let $w_{\leq t}=(w_{1},w_{2},\dots,w_{t})\in\mathcal{W}_{t}\coloneq\mathcal{V}^{t}$ be a sequence of tokens. We define a state and an action as $s_{t}\coloneq s_{0}\circ w_{<t}\in\mathcal{S}$ and $a_{t}\coloneq w_{t}\in\mathcal{A}$, where $\circ$ denotes concatenation and $s_{0}$ is a prompt (or a query). We will sometimes use $x$ and $y$ as substitute of $s$ and $a$ whenever required by context, and use the conventional prime notation $s^{\prime},a^{\prime}$ to denote $s_{t+1},a_{t+1}$ when only relative temporal index matters. The transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ is defined as simple concatenation $s^{\prime}=\mathcal{P}(s,a)\coloneq s\circ a$; $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is an unknown reward function; and $\gamma\in[0,1]$ is a discount factor.
@@ 2 Preliminaries
Soft value functions. The goal of soft RL is to maximize the expected return $J(\pi)=\mathbb{E}[\sum_{t}\gamma^{t}\tilde{r}_{t}\mid\pi,\mathcal{M}]$. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ and the state value function $V^{\pi}:\mathcal{S}\to\mathbb{R}$ are essentially recursive definition of the return $J(\pi)$. In particular, we consider soft value functions (entropy-regularized) [35, 16] in this work, which are defined as follows:
$Q^{\pi}(s,a)$ :-$\coloneq\textstyle\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}_{t}\mid\pi,\mathcal{M},S_{0}=s,A_{0}\!=\!a]={r}(s,a)+\gamma V^{\pi}(s^{\prime}),:$ (1)
$V^{\pi}(s)$ :-$\coloneq\mathbb{E}_{a\sim\pi(s)}[Q(s,a)-\log\pi(a\mid s)]:$ (2)
@@
In an operator notation, the soft Bellman operator $\mathcal{B}^{\pi}_{r}:\mathbb{R}^{\mathcal{S}\times\mathcal{A}}\to\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ is defined as $(\mathcal{B}^{\pi}_{r}Q)(s,a)=r(s,a)+\gamma V^{\pi}(s^{\prime})$. Note this soft Bellman (evaluation) operator is a contractor [36], in contrast to the non-contractive soft Bellman improvement operator [35]. We will throughout this paper use soft Bellman operator to denote $\mathcal{B}^{\pi}_{r}$, the policy evaluation operator.
Remark on $\mathcal{B}^{\pi}_{r}$. The contractive property of $\mathcal{B}^{\pi}_{r}$ defines a unique soft $Q$-function, for a reward function $r$, as the fixed point of $Q=\mathcal{B}^{\pi}_{r}Q$, given the Banach fixed-point theorem [37]. The contraction property of $\mathcal{B}^{\pi}_{r}$ later helps us to obtain our results in Section 3.
@@
Max-entropy IRL. With these definitions, we are now ready to introduce our choice of concrete IRL algorithm, IQL [38], which is an extension of maximum-entropy IRL [16]. We start with the Generative Adversarial IL (GAIL) [17] objective function:
@@
Inverse soft Q-learning. IQL [38] is a recent popular extension of GAIL, which is also the choice of algorithm of our paper. It’s objective function is defined as:
$\textstyle\max_{Q}\min_{\pi}\mathcal{J}(\pi,Q)$ :-$\coloneq\mathbb{E}_{\rho^{\star}}[Q(s,a)-\gamma V^{\pi}(s^{\prime})]-(1-\gamma)\mathbb{E}_{s_{0}}[V^{\pi}(s_{0})]-\psi(\mathcal{T}^{\pi}Q),:$ (4)
$\textstyle\Leftrightarrow\quad\max_{Q}\mathcal{J}^{\star}(Q)$ :-$\coloneq\mathbb{E}_{\rho^{\star}}[\phi(Q(s,a)-\gamma V^{Q}(s^{\prime}))]-(1-\gamma)\mathbb{E}_{s_{0}}[V^{Q}(s_{0})]:$ (5)
@@
The core idea is can be summarized as: (i) The original GAIL objective $\min_{\pi}\max_{r}L(\pi,r)$ can be equivalently swapped as $\max_{r}\min_{\pi}L(\pi,r)$ by strong duality, given proper regularization and feasible domains; (ii) For a fixed policy $\pi$, the saddle-point problem $\max_{r}\min_{\pi}L(\pi,r)$ for a fixed $\pi$ is provably equivalent to $\max_{Q}\min_{\pi}\mathcal{J}(\pi,Q)$. This equivalence allows direct optimization in the value-function space, eliminating the need to explicitly model the inverse reward $r$; (iii) For a fixed $Q$, it is well-known that the entropy-regularized policy has a closed-form solution given by $\pi_{Q}=\exp Q(s,a)/\sum_{a}\exp Q(s,a)$. And Danskin’s theorem [39] implies the saddle-point optimization $\max_{Q}\min_{\pi}\mathcal{J}(\pi,Q)$ can be casted as a simple maximization $\max_{Q}\mathcal{J}^{\star}(Q)$.
Remark on IQL. IQL reparameterizes the saddle-point formulation of IRL, enabling simpler and more stable optimization, and has thus recently gained popularity. We refer readers to [17, 38, 40], for a detailed derivation, background, and follow-up works. Throughout this paper, we use the deterministic form of IQL since the transition function $\mathcal{P}$ for LLMs is deterministic; the original stochastic formulation can be found in Garg et al. [38].
3 Method
3.1 Notation Setup
Defining different sort of backup operators with different desired properties has been a significant string of works in the RL literature. We refer to Asadi and Littman [41] and Miahi et al. [42] for detailed discussions. We start with the example of soft Bellman operator used in IQL
@@ 3.1 Notation Setup
$\mathcal{A}^{\star}_{p}(s):=\{a_{i}:\sum\nolimits_{i}\pi^{\star}(a_{i}\mid s)=p;~{}\pi^{\star}(a_{i}|s)\geq\pi^{\star}(a_{j}\mid s)~{}\forall~{}i,j\}\subseteq\mathcal{A}.$ (7)
where actions are sorted in descending order of probability density (i.e., $\pi^{\star}(a_{i}|s)\geq\pi^{\star}(a_{j}|s)$ for all $i<j$). Thus, for state $s$, the set $\mathcal{A}^{\star}_{p}(s)$ consists of actions with the largest probabilities whose cumulative probability equals $p$, a.k.a. top-$p$ tokens [43]. Although it’s uncommon to have a set whose cumulative probability equals exactly $p$, we use this notation for brevity, as the intended meaning is clear from context.
@@
Let $\pi^{\star}$ and $Q^{\star}$ be the optimal policy and its corresponding soft $Q$-function in $\mathcal{M}$, and let $\pi^{\star}_{p}$ and $\bar{Q}_{p}^{\star}$ denote their counterpart in $\mathcal{M}_{p}$. As $\mathcal{M}_{p}$ is also a well-defined MDP, the existence of $\pi^{\star},\pi^{\star}_{p},Q^{\star},\bar{Q}^{\star}_{p}$ follows directly from the fixed point theorem [37]. Note that we use $\pi^{\star}$ to refer to both optimal policy and the teacher model, as in IRL it is often assumed that the demonstrator (the teacher) is optimal.
@@
$\textsc{proj}_{p}\left(\pi\right)(a\mid s):=\begin{cases}\pi(a\mid s)/\sum_{a\in\mathcal{A}^{\star}_{p}}\pi(a\mid s),&\text{if~{}}a\in\mathcal{A}^{\star}_{p}\\ 0,&\text{otherwise}\end{cases}$ (8)
@@
$(\mathcal{B}^{\pi}_{p}\bar{Q})(s,a)=r(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textsc{proj}_{p}(\pi)}}[\bar{Q}(s^{\prime},a^{\prime})-\log{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textsc{proj}_{p}(\pi)}(a^{\prime}\mid s^{\prime})].$ (9)
@@ 4 Implementation
We choose IQL [38] as our base IRL algorithm, and show that how to practically implement its top-$p$ counterpart. Recall that the IQL objective is given by:
$\textstyle\max_{Q}\mathcal{J}^{\star}(Q)$ :-$\coloneq\mathbb{E}_{\rho^{\star}}[\phi(Q(s,a)-\gamma V^{\pi_{Q}}(s^{\prime}))]-(1-\gamma)\mathbb{E}_{s_{0}}[V^{\pi_{Q}}(s_{0})],\textstyle V^{\pi_{Q}}(s)\coloneq\log\sum_{a}\exp Q(s,a).:$ (13)
@@
     Sample mini-batches $\mathcal{D}^{\star}_{\text{mini}},\mathcal{D}^{PT}_{\text{mini}}$ from dataset $\mathcal{D}^{\star},\mathcal{D}^{PT}$ respectively.
     Compute $\mathcal{F}_{p}^{\star}$ according to Definition 7 and $Q(s,a)$ using the student logits.
     Compute $\pi_{Q}(a\mid s)=\exp Q(s,a)/\sum_{a}\exp Q(s,a)$ and $\textsc{proj}_{p}\pi_{Q}$ using Definition 3.
     Compute the value function using $V^{\textsc{proj}_{p}\pi_{Q}}(s)=\mathbb{E}_{\textsc{proj}_{p}\pi_{Q}}[Q(s,a)-\log\textsc{proj}_{p}\pi_{Q}(a\mid s)]$
     Compute $\mathcal{J}^{\star}$ in Eq. 16 and the language modeling objective $\mathcal{J}_{PT}$ for student model update.
@@
$(\mathcal{F}_{p}^{\star}Q)(s,a)\coloneq\begin{cases}Q(s,a),&\text{if~{}}a\in\mathcal{A}^{\star}_{p}\\ -\infty,&\text{otherwise}\end{cases}$ (14)
@@
Policy projection. We are also ready to handle the policy $\pi_{Q}$ through projection. By definition $\pi_{Q}(a\mid s)=\exp Q(s,a)/\sum_{a}\exp Q(s,a)$, as a result its projected counterpart is $(\textsc{proj}_{p}\pi_{Q})=\exp Q(s,a)/\sum_{\mathcal{A}^{\star}_{p}}\exp Q(s,a)$ for $a\in\mathcal{A}_{p}^{\star}$ otherwise 00.
@@
$\max_{Q}\mathcal{J}^{\star}(Q)\coloneq\mathbb{E}_{\rho^{\star}}[\phi((\mathcal{F}_{p}^{\star}Q)(s,a)-\gamma V^{\textsc{proj}_{p}\pi_{Q}}(s^{\prime}))]-(1-\gamma)\mathbb{E}_{s_{0}}[V^{\textsc{proj}_{p}\pi_{Q}}(s_{0})]$ (15)
Further IQL details. $\phi(\cdot)$ is a reward regularizer, we follow Garg et al. [38] to use $\chi^{2}$, corresponding to $\phi(x)=x-x^{2}/4\alpha$ for some coefficient $\alpha$. We use $\alpha=0.1$ for our experiments. In the IQL implementation, the second term in Eq. (13) is rewritten equivalently to a TD update through the telescopic identity [38]. This identity states: for any $\pi$, valid occupancy measure $\mu$, and value function $V^{\pi}$, we have $\mathbb{E}_{(s,a)\sim\mu}\left[V^{\pi}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime})\right]=(1-\gamma)\mathbb{E}_{s_{0}}[V^{\pi}(s_{0})].$
@@
$\max_{Q}\mathcal{J}^{\star}(Q)\coloneq\mathbb{E}_{\rho^{\star}}[\phi((\mathcal{F}_{p}^{\star}Q)(s,a)-\gamma V^{\textsc{proj}_{p}\pi_{Q}}(s^{\prime}))]-\mathbb{E}_{\rho^{\star}}\left[V^{\textsc{proj}_{p}\pi_{Q}}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{\textsc{proj}_{p}\pi_{Q}}(s^{\prime})\right]$ (16)
@@
Further implementation details. Following Gu et al. [19], we maximize a language modeling objective $\mathcal{J}_{PT}=\mathbb{E}_{(s,a)\sim\mathcal{D}^{PT}}\log\pi_{Q}(a\mid s)$ to retain performance on established NLP benchmarks. We refer to Algorithm 1 and Section 2.3 of Gu et al. [19] for further details. We clip the Q values using a minimum value $Q_{\text{min}}=-10$ for numerical stability.
5 Experiments
@@
Table 1: Main results: The Rouge-L scores, averaged over five random seeds, of our method and baselines across model families. Parameter sizes are listed at the top of each block. Best results per block are in bold. *: For GPT-2 class, the baseline results are directly duplicated from Gu et al. [19].

### v2

@@ Abstract
Introduction
The successes of large language models (LLMs) have brought transformative advancements across many NLP tasks, albeit often at the cost of substantial computational resources due to their massive size. Model distillation is a widely adopted approach to compressing these large, high-capacity models (teachers) into smaller, efficient models (students). This practice has become standard, ranging from proprietary ones (achiam2023gpt) to open-source model families (touvron2023llama; grattafiori2024llama; guo2025deepseek; bai2023qwen; yang2024qwen2).
The seminal work of knowledge distillation (KD) proposed by hinton2015distilling established the foundation for distribution-matching distillation methods, which were later extended to sequence/auto-regressive models (kim2016sequence). In distribution-matching methods, at each time step during text generation, a student model $\pi:\mathcal{X}\to\Delta(\mathcal{V})$ seeks to approximate the probability distribution, over an vocabulary $\mathcal{V}$, produced by the teacher $\pi^{\star}$, for example by minimizing a (forward) KL-divergence (sanh2019distilbert; jiao2019tinybert; wang2020minilm) $\min_{\pi}\mathbb{KL}[\pi^{\star}(x_{\leq t})\|\pi(x_{\leq t})]$, where $x_{\leq t}\in\mathcal{X}$ is a sequence of length $t$, and $\Delta(\Omega)$ denotes a probability simplex over space $\Omega$.
Notably, this sequential distribution matching can be interpreted as behavior cloning (BC) (pomerleau1991efficient), within the imitation learning (IL) and inverse reinforcement learning (IRL) literature (abbeel2004apprenticeship; russell1998learning; ng2000algorithms; ziebart2008maximum; ho2016generative). While there are nuanced differences between IL and IRL, we use them interchangeably.
@@ Introduction
Casting distillation as an IL problem is not new per se; indeed, many existing distillation works (kim2016sequence; agarwal2024policy; gu2023minillm; wen2023f) perform BC. Before reviewing these works through the lens of IL, we highlight two key distinctions that help characterize the distillation setup: (i) offline vs. online: This refers to whether distillation is performed on a fixed, pre-collected dataset (offline) or involves continuously generating new data during training (online). (ii) off-policy vs. mixed-policy: A method is considered purely on-policy if it uses only student-generated data for training; otherwise, it is categorized as off-policy. However, purely on-policy distillation is rare in practice—most methods use a mixture of teacher and student data. We refer to this setup as mixed-policy, though it is technically off-policy in the RL literature.
The pioneer work of sequence-level KD (SeqKD) (kim2016sequence) is an offline off-policy BC. Follow up BC style works fall into two main categories: (i) online mixed-policy (agarwal2024policy; gu2023minillm; ko2024distillm); and (ii) offline mixed-policy (wen2023f); potentially with different choice of divergence measure - forward/reverse KL (kim2016sequence; gu2023minillm), Jensen-Shannon divergence (agarwal2024policy), and general $f$-divergence (wen2023f).
Figure 1: We average the sorted token probabilities across 2020 sequences generated by Qwen-2.5 3B. Top 5050 tokens account for $96\%$ of the total mass, and the top $7$ tokens contribute $\geq 90\%$.
However, BC is known to suffer from compounding errors (ross2011reduction; tu2022sample), also referred as exposure bias (ranzato2015sequence) in the context of auto-regressive models. To address these issues, the IL literature offers a rich toolbox beyond BC, for example generative adversarial training (ho2016generative), occupancy matching (syed2008apprenticeship), and more. We refer readers to osa2018algorithmic for a comprehensive review. While not exhaustive, many modern IL methods (syed2008apprenticeship; ho2016generative; ziebart2008maximum) involve modeling an inverse reward function and then optimizing a policy with respect to it using RL. These RL algorithms are often based on temporal-difference (TD) learning (sutton1998reinforcement), which estimates the long-term impact of an action (or token) given a state (or sub-sequence) to mitigate the compounding errors.
Distillation may potentially benefit from the broader IL toolbox—particularly from TD learning methods. Earlier work has already demonstrated the potential of such approaches in smaller auto-regressive models. For instance, yu2017seqgan and wu2021textgail can be viewed as early applications of ho2016generative. For modern LLMs, jia2024adversarial explored TD-style IL methods for distillation, through value moment matching (abbeel2004apprenticeship; ziebart2008maximum; swamy2021moments) in particular. However, applying TD learning to LLMs is generally challenging due to their vast state-action space $\mathcal{V}^{T}$ (snell2022offline; yu2023mathcal; havrilla2023trlx).
@@
Preliminaries
MDP notations. We define the conventional MDP tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ as follows. Let $w_{\leq t}=(w_{1},w_{2},\dots,w_{t})\in\mathcal{W}_{t}\coloneq\mathcal{V}^{t}$ be a sequence of tokens. We define a state and an action as $s_{t}\coloneq s_{0}\circ w_{<t}\in\mathcal{S}$ and $a_{t}\coloneq w_{t}\in\mathcal{A}$, where $\circ$ denotes concatenation and $s_{0}$ is a prompt (or a query). We will sometimes use $x$ and $y$ as substitute of $s$ and $a$ whenever required by context, and use the conventional prime notation $s^{\prime},a^{\prime}$ to denote $s_{t+1},a_{t+1}$ when only relative temporal index matters. The transition function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ is defined as simple concatenation $s^{\prime}=\mathcal{P}(s,a)\coloneq s\circ a$; $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is an unknown reward function; and $\gamma\in[0,1]$ is a discount factor.
@@
Soft value functions. The goal of soft RL is to maximize the expected return $J(\pi)=\mathbb{E}[\sum_{t}\gamma^{t}\tilde{r}_{t}\mid\pi,\mathcal{M}]$. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ and the state value function $V^{\pi}:\mathcal{S}\to\mathbb{R}$ are essentially recursive definition of the return $J(\pi)$. In particular, we consider soft value functions (entropy-regularized) (littman1996algorithms; ziebart2008maximum) in this work, which are defined as follows:
$Q^{\pi}(s,a)$ $\coloneq\textstyle\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}\tilde{r}_{t}\mid\pi,\mathcal{M},S_{0}=s,A_{0}\!=\!a]$
$={r}(s,a)+\gamma V^{\pi}(s^{\prime}),$ (1)
$V^{\pi}(s)$ $\coloneq\mathbb{E}_{a\sim\pi(s)}[Q(s,a)-\log\pi(a\mid s)]$ (2)
@@
In an operator notation, the soft Bellman operator $\mathcal{B}^{\pi}_{r}:\mathbb{R}^{\mathcal{S}\times\mathcal{A}}\to\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ is defined as $(\mathcal{B}^{\pi}_{r}Q)(s,a)=r(s,a)+\gamma V^{\pi}(s^{\prime})$. Note this soft Bellman (evaluation) operator is a contractor (haarnoja2018soft), in contrast to the non-contractive soft Bellman improvement operator (littman1996algorithms). We will throughout this paper use soft Bellman operator to denote $\mathcal{B}^{\pi}_{r}$, the policy evaluation operator.
Remark on $\mathcal{B}^{\pi}_{r}$. The contractive property of $\mathcal{B}^{\pi}_{r}$ defines a unique soft $Q$-function, for a reward function $r$, as the fixed point of $Q=\mathcal{B}^{\pi}_{r}Q$, given the Banach fixed-point theorem (banach1922operations). The contraction property of $\mathcal{B}^{\pi}_{r}$ later helps us to obtain our results in Section Method.
@@
Max-entropy IRL. With these definitions, we are now ready to introduce our choice of concrete IRL algorithm, IQL (garg2021iq), which is an extension of maximum-entropy IRL (ziebart2008maximum). We start with the Generative Adversarial IL (GAIL) (ho2016generative) objective function:
@@
Inverse soft Q-learning. IQL (garg2021iq) is a recent popular extension of GAIL, which is also the choice of algorithm of our paper. It’s objective function is defined as:
$\textstyle\max_{Q}\min_{\pi}\mathcal{J}(\pi,Q)$ $\coloneq\mathbb{E}_{\rho^{\star}}[Q(s,a)-\gamma V^{\pi}(s^{\prime})]$
$-(1-\gamma)\mathbb{E}_{s_{0}}[V^{\pi}(s_{0})]-\psi(\mathcal{T}^{\pi}Q),$ (4)
$\textstyle\Leftrightarrow\quad\max_{Q}\mathcal{J}^{\star}(Q)$ $\coloneq\mathbb{E}_{\rho^{\star}}[\phi(Q(s,a)-\gamma V^{Q}(s^{\prime}))]$
$-(1-\gamma)\mathbb{E}_{s_{0}}[V^{Q}(s_{0})]$ (5)
@@
The core idea is can be summarized as: (i) The original GAIL objective $\min_{\pi}\max_{r}L(\pi,r)$ can be equivalently swapped as $\max_{r}\min_{\pi}L(\pi,r)$ by strong duality, given proper regularization and feasible domains; (ii) For a fixed policy $\pi$, the saddle-point problem $\max_{r}\min_{\pi}L(\pi,r)$ for a fixed $\pi$ is provably equivalent to $\max_{Q}\min_{\pi}\mathcal{J}(\pi,Q)$. This equivalence allows direct optimization in the value-function space, eliminating the need to explicitly model the inverse reward $r$; (iii) For a fixed $Q$, it is well-known that the entropy-regularized policy has a closed-form solution given by $\pi_{Q}=\exp Q(s,a)/\sum_{a}\exp Q(s,a)$. And Danskin’s theorem (danskin2012theory) implies the saddle-point optimization $\max_{Q}\min_{\pi}\mathcal{J}(\pi,Q)$ can be casted as a simple maximization $\max_{Q}\mathcal{J}^{\star}(Q)$.
Remark on IQL. IQL reparameterizes the saddle-point formulation of IRL, enabling simpler and more stable optimization, and has thus recently gained popularity. We refer readers to (ho2016generative; garg2021iq; al2023ls), for a detailed derivation, background, and follow-up works. Throughout this paper, we use the deterministic form of IQL since the transition function $\mathcal{P}$ for LLMs is deterministic; the original stochastic formulation can be found in garg2021iq.
Method
Notation Setup
Defining different sort of backup operators with different desired properties has been a significant string of works in the RL literature. We refer to asadi2017alternative and miahiresmax for detailed discussions. We start with the example of soft Bellman operator used in IQL
@@
$\textstyle\mathcal{A}^{\star}_{p}(s):=$
$\{a_{i}:\textstyle\sum_{i}\pi^{\star}(a_{i}\mid s)=p;~\pi^{\star}(a_{i}|s)\geq\pi^{\star}(a_{j}\mid s)~\forall~i,j\}\subseteq\mathcal{A}.$ (7)
where actions are sorted in descending order of probability density (i.e., $\pi^{\star}(a_{i}|s)\geq\pi^{\star}(a_{j}|s)$ for all $i<j$). Thus, for state $s$, the set $\mathcal{A}^{\star}_{p}(s)$ consists of actions with the largest probabilities whose cumulative probability equals $p$, a.k.a. top-$p$ tokens (holtzman2019curious). Although it’s uncommon to have a set whose cumulative probability equals exactly $p$, we use this notation for brevity, as the intended meaning is clear from context.
@@
Let $\pi^{\star}$ and $Q^{\star}$ be the optimal policy and its corresponding soft $Q$-function in $\mathcal{M}$, and let $\pi^{\star}_{p}$ and $\bar{Q}_{p}^{\star}$ denote their counterpart in $\mathcal{M}_{p}$. As $\mathcal{M}_{p}$ is also a well-defined MDP, the existence of $\pi^{\star},\pi^{\star}_{p},Q^{\star},\bar{Q}^{\star}_{p}$ follows directly from the fixed point theorem (banach1922operations). Note that we use $\pi^{\star}$ to refer to both optimal policy and the teacher model, as in IRL it is often assumed that the demonstrator (the teacher) is optimal.
@@
$\textsc{proj}_{p}\left(\pi\right)(a\mid s):=\begin{cases}\pi(a\mid s)/\sum_{a\in\mathcal{A}^{\star}_{p}}\pi(a\mid s),&\text{if }a\in\mathcal{A}^{\star}_{p}\\ 0,&\text{otherwise}\end{cases}$ (8)
@@
$(\mathcal{B}^{\pi}_{p}\bar{Q})(s,a)$ $=r(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textsc{proj}_{p}(\pi)}}[\bar{Q}(s^{\prime},a^{\prime})$ (9)
$-\log{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textsc{proj}_{p}(\pi)}(a^{\prime}\mid s^{\prime})].$ (10)
@@
We choose IQL (garg2021iq) as our base IRL algorithm, and show that how to practically implement its top-$p$ counterpart. Recall that the IQL objective is given by:
$\textstyle\max_{Q}\mathcal{J}^{\star}(Q)\coloneq\mathbb{E}_{\rho^{\star}}[\phi(Q(s,a)-\gamma V^{\pi_{Q}}(s^{\prime}))]$
$-(1-\gamma)\mathbb{E}_{s_{0}}[V^{\pi_{Q}}(s_{0})],\textstyle V^{\pi_{Q}}(s)\coloneq\log\sum_{a}\exp Q(s,a).$ (14)
@@
  Sample mini-batches $\mathcal{D}^{\star}_{\text{mini}},\mathcal{D}^{PT}_{\text{mini}}$ from datasets.
  Compute $\mathcal{F}_{p}^{\star}$ according to Definition 7 and $Q(s,a)$.
  Compute $\pi_{Q}(a\mid s)$ and $\textsc{proj}_{p}\pi_{Q}$ using Definition 3.
  Compute projected value function using $V^{\textsc{proj}_{p}\pi_{Q}}(s)$.
  Compute $\mathcal{J}^{\star}$ in Eq. Implementation and $\mathcal{J}_{PT}$ for student update.
@@
[…]
@@
Policy projection. We are also ready to handle the policy $\pi_{Q}$ through projection. By definition $\pi_{Q}(a\mid s)=\exp Q(s,a)/\sum_{a}\exp Q(s,a)$, as a result its projected counterpart is $(\textsc{proj}_{p}\pi_{Q})=\exp Q(s,a)/\sum_{\mathcal{A}^{\star}_{p}}\exp Q(s,a)$ for $a\in\mathcal{A}_{p}^{\star}$ otherwise 0.
@@
$\max_{Q}\mathcal{J}^{\star}(Q)$ $\coloneq\mathbb{E}_{\rho^{\star}}[\phi((\mathcal{F}_{p}^{\star}Q)(s,a)-\gamma V^{\textsc{proj}_{p}\pi_{Q}}(s^{\prime}))]$
$-(1-\gamma)\mathbb{E}_{s_{0}}[V^{\textsc{proj}_{p}\pi_{Q}}(s_{0})]$ (16)
Further IQL details. $\phi(\cdot)$ is a reward regularizer, we follow garg2021iq to use $\chi^{2}$, corresponding to $\phi(x)=x-x^{2}/4\alpha$ for some coefficient $\alpha$. We use $\alpha=0.1$ for our experiments. In the IQL implementation, the second term in Eq. (14) is rewritten equivalently to a TD update through the telescopic identity (garg2021iq). This identity states: for any $\pi$, valid occupancy measure $\mu$, and value function $V^{\pi}$, we have $\mathbb{E}_{(s,a)\sim\mu}\left[V^{\pi}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{\pi}(s^{\prime})\right]=(1-\gamma)\mathbb{E}_{s_{0}}[V^{\pi}(s_{0})].$
@@
$\max_{Q}\mathcal{J}^{\star}(Q)$ $\coloneq\mathbb{E}_{\rho^{\star}}[\phi((\mathcal{F}_{p}^{\star}Q)(s,a)-\gamma V^{\textsc{proj}_{p}\pi_{Q}}(s^{\prime}))]$
$-\mathbb{E}_{\rho^{\star}}\left[V^{\textsc{proj}_{p}\pi_{Q}}(s)-\gamma\mathbb{E}_{s^{\prime}}V^{\textsc{proj}_{p}\pi_{Q}}(s^{\prime})\right]$ (17)
@@
Further implementation details. Following gu2023minillm, we maximize a language modeling objective $\mathcal{J}_{PT}=\mathbb{E}_{(s,a)\sim\mathcal{D}^{PT}}\log\pi_{Q}(a\mid s)$ to retain performance on established NLP benchmarks. We refer to Algorithm 1 and Section 2.3 of gu2023minillm for further details. We clip the Q values using a minimum value $Q_{\text{min}}=-10$ for numerical stability.
Experiments
@@
Figure 2: Comparison of win rates against KD, SeqKD, and MiniLLM baselines. We evaluate using GPT-4o-mini (openai2024gpt4omini) as the judging oracle, with Qwen-2.5 (3B) as the teacher model and a smaller (0.5B) model as the student. Results are based on 500 responses per distilled model, generated under the Dolly evaluation setting.
Table 1: Main results: The Rouge-L scores, averaged over five random seeds, of our method and baselines across model families. Parameter sizes are listed at the top of each block. Best results per block are in bold. *: For GPT-2 class, the baseline results are directly duplicated from gu2023minillm.